OpenStudy (anonymous):

The probability, P, that a person responds to an advertisement can be modeled by the exponential function P = 1 – e^-0.047t. In the formula, t is the number of days since the advertisement first appeared. How many days will it take for the probability that a person has responded reach 75% ?

OpenStudy (anonymous):

0.75 = 1 - exp(-0.047 t) exp(-0.047 t) = 1-0.75 = 0.25 -0.047 t = ln(0.25) = -1.386 solve for t